Circular No. 6824

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

QX NORMAE S. Corbel, Centre d’Etudes Atomiques, Saclay; D. Hannikainen, Uni-

versity of Helsinki Observatory; and J. F. Gonzalez and O. R. Hainaut, European Southern Observatory (ESO), report: “Following the recent xray outburst of the soft x-ray transient 4U 1608–52 detected by RXTE (IAUC 6819), observations with the ESO 3.5-m New Technology Telescope at La Silla revealed that the optical counterpart QX Nor is also flaring. Total magnitudes on Feb. 16 (eight days after soft-x-ray maximum) are V = 20.07 ± 0.06, R = 18.92 ± 0.04, and I = 17.78 ± 0.05, which indicate a brightening of more than 3 mag in R, compared to the quiescent state (Wachter 1997, Ap.J. 485, 839). Radio observations and optical spectroscopy, as well as follow-up photometry, are strongly encouraged.”

COMET 55P/TEMPEL-TUTTLE T. L. Hayward, Cornell University; and M. S. Hanner, Jet Propulsion

Laboratory, California Institute of Technology, write: “We observed comet 55P in the mid-infrared on Jan. 14.5 UT with the 5-m Hale telescope (+ SpectroCam-10). Imaging at 10.3 μm revealed a point-like source with flux density 2.1 ± 0.2 Jy. A spectrogram (range 8–13 μm; resolution 100) showed that the emission was close to a 320-K blackbody, compared to an equilibrium temperature of 253 K for r = 1.215 AU. A weak silicate feature may be present ∼ 5 percent above the continuum. If the emission were from a bare nucleus at 320 K, then the observed flux implies a nuclear radius of 2 km.”

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: Jan. 26.60 UT, 8.1 (A. Pearce, Ned-

lands, W. Australia, 20×80 binoculars); Feb. 1.03, 8.2 (J. G. de S. Aguiar, Campinas, Brazil, 0.23-m reflector); 7.60, 8.4 (Pearce); 16.50, 8.6 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); 18.61, 8.4 (Pearce).

COMET 29P/SCHWASSMANN-WACHMANN 1 Visual m_{1} estimates: Feb. 1.06 UT, 13.1 (K. Sárneczky, Ráktanya,

Hungary, 0.25-m reflector); 5.17, 12.5: (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

Corrigendum. On IAUC 6816, lines 1–2, for Jan. 14.25 and 24.25 UT read Jan. 15.25 and 25.25 UT

1998 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green